OpenStudy (anonymous):

decide whether the equation is a trigonometric identity. cos^2 (x)(1+tan^2 (x))=1 sec(x)tan(x)(1-sin^2(x)=sin(x)

OpenStudy (anonymous):

I will fan and medal

OpenStudy (zzr0ck3r):

\(\cos^2(x)(1+\tan^2(x))=\cos^2(x)+\cos^2(x)\dfrac{\sin^2(x)}{\cos^2(x)}=\cos^2(x)+\sin^2(x)=?\)

OpenStudy (dinamix):

=1

OpenStudy (zzr0ck3r):

For the second one, \[\cos^2(x)=1-\sin^2(x)\] So you have \[\sec(x)\tan(x)\cos^2(x)=\frac{1}{\cos(x)}\frac{\sin(x)}{\cos(x)}\cos^2(x)=?\]

OpenStudy (dinamix):

\[\sec(x) = \frac{ 1 }{ \cos(x)}\] i help u litttle mate

OpenStudy (dinamix):

@zzr0ck3r why give him answer directly ?

OpenStudy (dinamix):

thats not good for him

OpenStudy (dinamix):

right ?

OpenStudy (anonymous):

thanks both of you for the help